A monochromatic light ray with the wavelengths of 600nm disperses in glass with a refractive index of 1.5. Find the wavelength of the light ray in the glass


1
Expert's answer
2021-12-19T18:17:03-0500

We know that

"\lambda=600nm\\\mu=1,\mu'=1.5"

"\lambda\alpha\frac{1}{\mu}"

Now

"\frac{\lambda}{\lambda'}=\frac{\mu'}{\mu}"

"\lambda'=\frac{\mu}{\mu'}\times\lambda"

"\lambda'=600\times10^{-9}\times\frac{1}{1.5}=400nm"